OpenStudy (anonymous):

A high school basketball team has a budget specifically for towels and extra basketballs. A towel costs $6 and a basketball costs $30. Assume the budget is $300 and that 25 towels are needed for the team for the entire season. Use a standard form equation to determine how many basketballs may be purchased by the team.

OpenStudy (anonymous):

300=6t+30b

OpenStudy (anonymous):

25 towels is $150.

OpenStudy (anonymous):

is that what you need?

OpenStudy (anonymous):

150/30 = 5

OpenStudy (anonymous):

So 5 basketballs can be purchased.

OpenStudy (anonymous):

i need this answer:Use a standard form equation to determine how many basketballs may be purchased by the team.

OpenStudy (anonymous):

30b+6t=300

OpenStudy (anonymous):

There's your answer.

OpenStudy (anonymous):

here are the options: A. 5 B. 15 C. 30 D. 275

OpenStudy (anonymous):

I said it's 5.

OpenStudy (anonymous):

So it's A.

OpenStudy (anonymous):

This is a really easy question.

OpenStudy (anonymous):

Rewrite the equation below into standard form.

OpenStudy (anonymous):

Ughh.

OpenStudy (anonymous):

25t+5b=300.

OpenStudy (anonymous):

-(3/2)x=-(2/5)y-(10/3) Rewrite the equation below into standard form. 1 2 3 4 5 6 7 8 9 10 Time Remaining 44:01 Rewrite the equation below into standard form. A. B. C. D. Please select the best answer from the choices provided A B C D Mark this and returnNextSubmit

OpenStudy (anonymous):

i need this answerd: -(3/2)x=-(2/5)y-(10/3) Rewrite the equation below into standard form.